# Token Preference Optimization with Self-Calibrated Visual-Anchored Rewards for Hallucination Mitigation

### v3

@@ Abstract
Direct Preference Optimization (DPO) has been demonstrated to be highly effective in mitigating hallucinations in Large Vision Language Models (LVLMs) by aligning their outputs more closely with human preferences. Despite the recent progress, existing methods suffer from two drawbacks: 1) Lack of scalable token-level rewards; and 2) Neglect of visual-anchored tokens. To this end, we propose a novel Token Preference Optimization model with self-calibrated rewards (dubbed as TPO), which adaptively attends to visual-correlated tokens without fine-grained annotations. Specifically, we introduce a token-level visual-anchored reward as the difference of the logistic distributions of generated tokens conditioned on the raw image and the corrupted one. In addition, to highlight the informative visual-anchored tokens, a visual-aware training objective is proposed to enhance more accurate token-level optimization. Extensive experimental results have manifested the state-of-the-art performance of the proposed TPO. For example, by building on top of LLaVA and Qwen, our TPO boosts the performance absolute improvement for hallucination benchmarks.
@@
Jihao Guthanks: Equal contribution.thanks: Work done during an internship at Alibaba Group.1, Yingyao Wang11footnotemark: 11, Meng Cao2, Pi Bu1, Jun Songthanks: Corresponding Author.1, Yancheng He1, Shilong Li1, Bo Zheng1 1Alibaba Group 2Mohamed bin Zayed University of Artificial Intelligence {gujihao.gjh, wangyingyao.wyy, jsong.sj}@taobao.com
@@ 1 Introduction
Methods Visual- Anchored Token- level Non Fine-grained Annotations
@@
Table 1: Comparisons with hallucination mitigation methods from the perspective of whether attending to vision-anchored tokens, whether generating token-level rewards and whether requiring fine-grained annotations. The compared methods include DPO Rafailov et al. (2024), POVID Zhou et al. (2024a), CSR Zhou et al. (2024b), RLHF-V Yu et al. (2024b), V-DPO Xie et al. (2024) and our proposed TPO.
@@ 2.1 LVLMs’ Hallucination
Previous studies have mitigated hallucinations by enhancing training data quality, refining decoding strategies, and post-processing generated responses Huang et al. (2024); Leng et al. (2024); Yu et al. (2024a); Han et al. (2024); Chen et al. (2024); Zhou et al. (2023); Yin et al. (2023); Lee et al. (2023); Shao et al. (2024). While these methods can lead to more accurate responses, they do not fundamentally resolve the issue of inadequate visual information association in LVLMs.
@@ 2.2 Preference Learning Methods
Several studies based on DPO focus on developing more robustly constructed preference data. For example, the POVID Zhou et al. (2024a) method constructs negative samples for preferred data by adding noise to the image and providing hallucinated patterns to guide the model to generate hallucinated responses. The pre-structured data will be employed for off-policy DPO training. Apart from these works, RLAIF Yu et al. (2024c) and CSR Zhou et al. (2024b) methods, which are built upon on-policy DPO strategy, construct preference pairs by iteratively performing self-rewarding to select preference pairs. A notable commonality among the aforementioned studies is that assigning response-level rewards for each generated sequence is insufficient for effectively aligning with genuinely hallucination-prone contents.
Other studies, RLHF-V Yu et al. (2024b) and V-DPO Xie et al. (2024), investigated this issue and achieved more fine-grained alignment of preference data. Nevertheless, this approach depends on resource-intensive annotations or data constructions and applies a fixed reward to all hallucinated segments, thus failing to account for the differing levels of relevance these segments may have to visual information. It is worth mentioning that CSR also considered this problem and introduced CLIP Radford et al. (2021) to calculate the relevance score between generated text and vision information as an additional reward. However, this method requires the introduction of an additional model, which reduces the training efficiency.
@@ 3.2 Visual-Anchored Rewards
where $\xi$ is a predefined noise parameter derived from a list with 1,000 equally spaced elements [^1] More details can be found in Appendix A, and experimental analysis can be found in Appendix E. $\bar{\xi}_{k}$ is a cumulant, i.e., $\bar{\xi}_{k}=\prod_{i=0}^{k}\xi_{i}$.
@@
where $s_{y_{i}}$ denotes the distribution difference of the token $y_{i}$ of the response $y$. $p_{log}$ refers to the raw logits output of the model, before applying softmax normalization. One example case is demonstrated in Figure 1, which demonstrates that $s$ reflects the visual relevance of each token $y_{i}$.
@@ 3.3 Token Preference Optimization
Compared to the original reward function in DPO (Eq (1)), we multiply each $p(y_{i}|x,v,y_{<i})$ by the generated visual-anchored rewards $c_{y_{i}}$ at the token level. $c_{y_{i}}^{\theta}$ is continuously updated at each step during training as the model changes. To calculate each token in the entire reward function, we add a term $log\frac{c_{y_{i}}^{\theta}}{c_{y_{i}}^{\text{ref}}}\in(-log3,log3)$, which has a reasonable upper and lower bound. For positive samples, this term is expected to increase, while for negative samples, it is expected to decrease. Due to the different methods of calculating $c_{y_{i}}$ that we set in Eq (5), this will encourage the increase of $s_{y_{i}}$ during the training process, making the token generation focus more on visual information.
@@
Acc F1 Score Hal $\downarrow$ Easy Hard aAcc SEED MMB LLaVA MM-Vet
@@
Table 2: Performence of LLaVA-1.5 on hallucination and general benchmarks. Score and Hall refer to the overall GPT-4 Achiam et al. (2023) score and hallucination rate, respectively. Easy represents the accuracy of with original images, hard represents the accuracy with manually edited challenging images, and aAcc is the average accuracy for each question. The results for POVID Zhou et al. (2024a) and CSR Zhou et al. (2024b) are based on our testing of their open-source model weights, while the results for V-DPO Xie et al. (2024) are taken from previous work (bold: the best score; underline: the second best)
Method AMBER MMHal HallusionBench
[…]
Qwen2-vl 86.5 90.0 3.5 29.0 67.0 48.8 64.0
+DPO 86.5 90.0 3.7 28.1 67.3 49.3 64.5
+TPO 86.4 89.9 4.2 18.8 67.9 50.0 65.2
Table 3: Performence of Qwen2-vl-7B on hallucination benchmarks.
@@ 4.1 Setup
Aligning with previous DPO-based approaches on hallucination mitigation, we mainly adopt the popular LVLM, LLaVA-1.5 Liu et al. (2024b), as the backbone model to validate the effectiveness of our TPO. Furthermore, to evaluate the effectiveness of TPO on more advanced and powerful model, we implement TPO training based on Qwen2-vl Wang et al. (2024), and compare it with the DPO method. For the dataset, we directly utilize the preference pairs provided by RLHF-V (5K) without their fine-grained human annotations.
@@
We primarily conduct the experiments on three hallucination benchmarks: AMBER Wang et al. (2023), MMHal-Bench Sun et al. (2023), and HallusionBench Guan et al. (2024). In this section, we mainly focus on AMBER’s discriminative task and report the accuracy and F1 metrics referencing Yu et al. (2024c). In addition, we provide the results of its Chair metric in Appendix D. Moreover, we also evaluate the performance of TPO on four general benchmarks: SEED Bench Li et al. (2023a), MMBench Liu et al. (2025), LLaVA Bench Liu et al. (2024c) and MM-Vet Yu et al. (2023). These benchmarks are used to evaluate the performance of the models on general tasks after hallucination alignment.
@@
We mainly compare TPO with the LLaVA-1.5-7B SFT model, as well as with the DPO and V-DPO Xie et al. (2024) methods trained using RLHF-V Yu et al. (2024b) data, along with two improved methods, CSR Zhou et al. (2024b) and POVID Zhou et al. (2024a). Moreover, to evaluate the effectiveness and robustness of TPO as the model size increases, we further evaluate the performance of TPO on the LLaVA-1.5-13B model and compared it with DPO. Additionally, to demonstrate the advantages of TPO, we reproduced the strong baseline method, RLHF-V, on LLaVA-1.5-13B and conducted a comparison.
To further demonstrate the effectiveness of TPO on more advanced and powerful model, we additionally employ Qwen2-vl-7B as the baseline mode and compare our TPO with DPO.
@@
Acc F1 Score Hal $\downarrow$ Easy Hard aAcc SEED MMB LLaVA MM-Vet
@@
4.2 Main Results
[…]
Notably, on the HallusionBench evaluation metrics, "Easy" represents the accuracy of original image-based questions, which tend to rely on prior knowledge, while "Hard" represents the accuracy of questions based on manually edited images, which tend to rely on visual information. Our method leads to the most significant improvement for the original model on hard questions, with a slight improvement on easy questions. This indicates that, compared to other methods, our approach enables the model to focus more on visual information rather than the textual prior knowledge to provide accurate answers.
On general benchmarks, our approach remains stable against the original LLaVA model and achieves the greatest improvement on most benchmarks. We attribute it to that our method helps the model associate with more visual information when answering questions. This shows that our approach can improve hallucination issues while maintaining good performance on general evaluation tasks. The results also indicate that TPO introduces performance fluctuations on the LLaVABench and MMVet metrics for LLaVA-1.5-13B. This occurs because the training data primarily focuses on aligning preferences within the context of hallucinations, which inevitably leads to slight overfitting. Although augmenting the dataset with task-specific data could mitigate this issue, it is not the research focus of this paper.
4.3 Results on Qwen2-vl
As Table 3 shown, we report the results on the key metrics of three hallucination benchmarks.
The results indicate that our TPO outperforms DPO on most benchmarks. On Qwen2-vl, which has strong inherent capabilities, using 5K RLHF-V data for DPO alignment barely improves the performance. However, introducing TPO leads to a significant further enhancement. This not only demonstrates the effectiveness of TPO on other powerful models but also shows that TPO’s token-level visual anchor rewards brings higher data utilization efficiency. Although Qwen2-vl shows almost no room for improvement on AMBER, TPO still achieves significant improvement on the Chair metric, as shown on Figure 5. TPO can capture and learn more subtle preferences from the data.
4.4 Ablation Studies
Visual-Anchored Rewards
Hyperparameters
To optimize the hyperparameters in TPO, we perform comparative experiments on the noise steps (Section 3.2) and parameter $a$ in Equation 5. As the Figure 5 shown, the model performs best with 500 noise steps. Testing $a$ across the range $a=[0,0.5,1]$, we find that $a$=0.5 produces the best outcome. This supports our hypothesis that setting $s=0$ and $c=1$, without adding extra reward signals, leads to superior performance [^2] The more detail results and discussions are provided in Appendix..
Figure 5: Performance curves with the change of the noise steps-(a) and the change of parameter $a$-(b), We separately present the F1 of AMBER, the hallucination rate of MMHAL, the aACC of HallusionBench, and the average value of the general benchmarks. More detailed metrics can be found in the Appendix B.
@@ 5 Conclusion
In this study, we propose a novel pereference alignment method, TPO, to mitigation hallucinations in LVLMs. TPO incorporates a self-calibrated visual-anchored reward mechanism that automatically identifies "vision-anchored tokens" and adaptively assigns appropriate rewards to them. By adding noise to the visual input and capturing changes in the generation probability of each token, TPO computes a score indicating each token’s relevance to visual information. Subsequently, a self-calibration process adjusts these scores to ensure that: 1) Rewards for positive examples exceed those for negative examples. 2) The final rewards enhance the focus on visual information for all of "vision-anchored tokens" in preference data. Based on the self-calibrated visual-anchored reward, TPO can perform more efficient token-level preference alignment optimization for LVLMs. Experimental results have proved that TPO not only alleviates the hallucination problem but also strengthens the model’s attention to visual input when generating responses.
@@ 6 Limitation
Although our method has achieved outstanding performance in addressing the hallucination problem, the self-calibrated visual-anchored rewards approach we used in this paper can be extended to even broader areas. By altering the way noise is added to images, we can shift from adding noise to the entire image to adding noise to specific key objects. It can enable the model to specifically improve its focus on image information in certain domains, thus having extensive industrial applications. We will continue to expand in this direction, and we believe that the technology we have proposed in this paper has a vast space for further development and application.
@@ A.1 Setup
We also trained the Qwen2-vl model. For our TPO method and the vanilla DPO method, we set the maximum learning rate to 5e-9 and trained for 4 epochs.
@@
Our experiments were all conducted on a server equipped with 8 Nivdia A100 GPUs; in specific cases (such as the 13B model), we utilized 32 Nivdia A100 GPUs. For the hyperparameter settings, all hyperparameters are consistent with those of our main experiment. Moreover, the level of diffusion noise in our model is represented by a formula $\xi=Sigmoid(l_{t})\times(0.5\times 10^{-2}-10^{-5})+10^{-5}$, where $l_{t}$ is a list of 1,000 numbers taken at equal intervals over the interval $[-6,6]$, and $\epsilon\in N(0,1)$.

### v4

@@ Abstract
Direct Preference Optimization (DPO) has been demonstrated to be highly effective in mitigating hallucinations in Large Vision Language Models (LVLMs) by aligning their outputs more closely with human preferences. Despite the recent progress, existing methods suffer from two drawbacks: 1) Lack of scalable token-level rewards; and 2) Neglect of visual-anchored tokens. To this end, we propose a novel Token Preference Optimization model with self-calibrated rewards (dubbed as TPO), which adaptively attends to visual-correlated tokens without fine-grained annotations. Specifically, we introduce a token-level visual-anchored reward as the difference of the logistic distributions of generated tokens conditioned on the raw image and the corrupted one. In addition, to highlight the informative visual-anchored tokens, a visual-aware training objective is proposed to enhance more accurate token-level optimization. Extensive experimental results have manifested the state-of-the-art performance of the proposed TPO. For example, by building on top of LLaVA and Qwen, our TPO boosts the performance absolute improvement for hallucination benchmarks. Our code is available at https://github.com/alibaba/TPO.
@@
Jihao Guthanks: Equal contribution.thanks: Work done during an internship at Alibaba Group.1, Yingyao Wang11footnotemark: 11, Meng Cao2, Pi Bu1, Jun Songthanks: Corresponding Author.1, Yancheng He1, Shilong Li1, Bo Zheng1 1Taobao & Tmall Group of Alibaba 2Mohamed bin Zayed University of Artificial Intelligence {gujihao.gjh, wangyingyao.wyy, jsong.sj}@taobao.com
@@ 1 Introduction
Methods
Visual-
Anchored
Token-
level
Non Fine-grained
Annotations
@@ Methods
Table 1: Comparisons with hallucination mitigation methods from the perspective of whether attending to vision-anchored tokens, whether generating token-level rewards and whether requiring fine-grained annotations. The compared methods include DPO Rafailov et al. (2024), POVID Zhou et al. (2024a), CSR Zhou et al. (2024b), MDPO Wang et al. (2024a), V-DPO Xie et al. (2024), RLHF-V Yu et al. (2024b).
@@ 2.1 LVLMs’ Hallucination
Previous studies have mitigated hallucinations by enhancing training data quality, refining decoding strategies, and post-processing generated responses Huang et al. (2024); Leng et al. (2024); Yu et al. (2024a); Han et al. (2024); Chen et al. (2024); Zhou et al. (2023); Yin et al. (2023); Lee et al. (2023); Shao et al. (2024a); Jiang et al. (2024); Yue et al. (2024); Xiao et al. (2025); Sarkar et al. (2024); Zhao et al. (2023). While these methods can lead to more accurate responses, they do not fundamentally resolve the issue of inadequate visual information association in LVLMs.
@@ 2.2 Preference Learning Methods
Several studies based on DPO focus on developing more robustly constructed preference data. For example, the POVID Zhou et al. (2024a) method constructs negative samples for preferred data by adding noise to the image and providing hallucinated patterns to guide the model to generate hallucinated responses. The MDPO incorporate optimization for image preference, training with the images before and after alteration as positive and negative samples. Apart from these works, RLAIF Yu et al. (2024c) and CSR Zhou et al. (2024b) methods, which are built upon on-policy DPO strategy, construct preference pairs by iteratively performing self-rewarding to select preference pairs. R1-Onevision Yang et al. (2025) enhances the visual reasoning capabilities by employing Group Relative Policy Optimization (GRPO). However, assigning response-level rewards for each generated sequence is insufficient for effectively aligning with genuinely hallucination-prone contents.
Other studies, RLHF-V Yu et al. (2024b) and V-DPO Xie et al. (2024), investigated this issue and achieved more fine-grained alignment of preference data. Nevertheless, this approach depends on resource-intensive annotations or data constructions and applies a fixed reward to all hallucinated segments, thus failing to account for the differing levels of relevance these segments may have to visual information. It is worth mentioning that CSR also considered this problem and introduced CLIP Radford et al. (2021) to calculate the relevance score between generated text and vision information as an additional reward, and TLDR Fu et al. (2024) score each token by training a scoring model. However, these methods requires the introduction of an additional model, which reduces the training efficiency.
@@ 3.2 Visual-Anchored Rewards
where $\xi$ is a predefined noise parameter derived from a list with 1,000 equally spaced elements [^1] More details can be found in Appendix A, and experimental analysis can be found in Appendix D. $\bar{\xi}_{k}$ is a cumulant, i.e., $\bar{\xi}_{k}=\prod_{i=0}^{k}\xi_{i}$.
@@
where $s_{y_{i}}$ denotes the distribution difference of the token $y_{i}$ of the response $y$. $p_{log}$ refers to the raw logits output of the model, before applying softmax normalization.
The sole variation in the input when computing $s_{y_{i}}$ of the token lies in whether noise is introduced into the image. A higher contrastive score, reflecting a greater divergence in logits, indicates that the current token is more susceptible to image perturbations. This increased sensitivity suggests a stronger dependence on and relevance to visual information, thereby marking it as a visual-anchored token. One example case is demonstrated in Figure 1, which demonstrates that $s_{y_{i}}$ reflects the visual relevance of each token $y_{i}$.
@@ 3.3 Token Preference Optimization
Compared to the original reward function in DPO (Eq (1)), we multiply each $p(y_{i}|x,v,y_{<i})$ by the generated visual-anchored rewards $c_{y_{i}}$ at the token level. $c_{y_{i}}^{\theta}$ is continuously updated at each step during training as the model changes. To calculate each token in the entire reward function, we add a term $log\frac{c_{y_{i}}^{\theta}}{c_{y_{i}}^{\text{ref}}}\in(-log3,log3)$ (as we set $a=0.5$ in Equation (5)), which has a reasonable upper and lower bound. For positive samples, this term is expected to increase, while for negative samples, it is expected to decrease. Due to the different methods of calculating $c_{y_{i}}$ that we set in Eq (5), this will encourage the increase of $s_{y_{i}}$ during the training process, making the token generation focus more on visual information.
@@
Acc F1 Score Hal $\downarrow$ Easy Hard aAcc SEED MMB LLaVA MMVet
R1-Onevision 80.2 85.7 3.85 36.46 63.74 50.47 62.80 35.2 83.7 67.8
@@
Qwen2-VL-7B 86.5 90.0 3.5 29.0 67.0 48.8 64.0 45.0 79.0 82.4 61.4
+ DPO 86.5 90.0 3.7 28.1 67.3 49.3 64.5 45.0 79.0 81.9 60.2
+ TPO (Ours) 86.4 89.9 4.2 18.8 67.9 50.0 65.2 45.0 79.0 82.9 61.4
Table 2: Performence of LLaVA-1.5 on hallucination and general benchmarks. Score and Hall refer to the overall GPT-4 Achiam et al. (2023) score and hallucination rate, respectively. Easy represents the accuracy of with original images, hard represents the accuracy with manually edited challenging images, and aAcc is the average accuracy for each question. The results for POVID Zhou et al. (2024a) and CSR Zhou et al. (2024b) are based on our testing of their open-source model weights, while the results for V-DPO Xie et al. (2024), MDPO Wang et al. (2024a) are taken from previous work
@@ 4.1 Setup
Aligning with previous DPO-based approaches on hallucination mitigation, we mainly adopt the popular LVLM, LLaVA-1.5 Liu et al. (2024b), as the backbone model to validate the effectiveness of our TPO. Furthermore, to evaluate the effectiveness of TPO on more advanced and powerful model, we implement TPO training based on Qwen2-VL Wang et al. (2024b), and compare it with the DPO method. For the dataset, we directly utilize the preference pairs provided by RLHF-V (5K) without their fine-grained human annotations.
@@
We primarily conduct the experiments on three hallucination benchmarks: AMBER Wang et al. (2023), MMHal-Bench Sun et al. (2023), and HallusionBench Guan et al. (2024). In this section, we mainly focus on AMBER’s discriminative task and report the accuracy and F1 metrics referencing Yu et al. (2024c). In addition, we provide the results of its Chair metric in Appendix C. Moreover, we also evaluate the performance of TPO on four general benchmarks: SEED Bench Li et al. (2023a), MMBench Liu et al. (2025), LLaVA Bench Liu et al. (2024c) and MM-Vet Yu et al. (2023). These benchmarks are used to evaluate the performance of the models on general tasks after hallucination alignment.
@@
We mainly compare TPO with the R1-Onevision Yang et al. (2025), LLaVA-1.5-7B SFT model, as well as with the V-DPO Xie et al. (2024) and DPO methods trained using RLHF-V Yu et al. (2024b) data, along with two improved methods, CSR Zhou et al. (2024b) and POVID Zhou et al. (2024a). Moreover, to evaluate the effectiveness and robustness of TPO as the model size increases, we further evaluate the performance of TPO on the LLaVA-1.5-13B model and compared it with DPO. Additionally, to demonstrate the advantages of TPO, we reproduced the strong baseline method, RLHF-V, on LLaVA-1.5-13B and conducted a comparison. Furthermore, we additionally employ Qwen2-VL-7B Wang et al. (2024b) as the baseline mode and compare our TPO with DPO.
4.2 Main Results
[…]
Notably, on the HallusionBench evaluation metrics, "Easy" represents the accuracy of original image-based questions, which tend to rely on prior knowledge, while "Hard" represents the accuracy of questions based on manually edited images, which tend to rely on visual information. Our method leads to the most significant improvement for the original model on hard questions. This indicates that our approach enables the model to focus more on visual information rather than textual prior knowledge to provide accurate answers.
In general benchmarks, our approach remains stable against the backbone models and achieves improvement on most benchmarks. We attribute it to our method helping the model associate with more visual information when answering questions. This shows that our approach can improve hallucination issues while maintaining good performance in general evaluation tasks.
4.3 Results on Qwen2-VL-7B
As Table 2 shown, we report the results on the key metrics of three hallucination benchmarks. The results indicate that our TPO outperforms DPO on most benchmarks. On Qwen2-VL-7B, which has strong inherent capabilities, using 5K RLHF-V data for DPO alignment barely improves the performance. However, introducing TPO leads to a significant further enhancement. This demonstrates that TPO can capture and learn more subtle preferences from the data and brings higher data utilization efficiency. In addition, the results on the chair metric in Figure 6 further demonstrate that TPO can also significantly solve the object hallucination problem of Qwen2-VL-7B.
4.4 Ablation Studies
Visual-Anchored Rewards
Noise Step
We conduct an ablation study on the noise steps, as illustrated in Figure 5 (a) and detailed in Table 5 of Appendix. The results indicate that optimal performance is achieved at 500 steps. This intermediate level of corruption allows the model to retain the general structure of the image while obscuring fine-grained details, thereby reducing the tendency to generate hallucinations among visual-anchored tokens.
These findings are consistent with previous work Zhou et al. (2024a) which noted that as image noise increases, models tend to produce responses such as “there are noise spots in the image,” while the probability of hallucination first increases and then decreases. Our experimental results align with this trend and provide further empirical support.
Parameter $\bf a$
We conduct experiments by varying the parameter $a$ introduced in Equation (5) with the results shown in Figure 5 (b) (detailed in Table 6 of Appendix). By setting $a=[0,0.5,1]$, we observed consistently good performance across all configurations. This suggests that effective performance is achieved as long as the reward mechanism successfully highlights token differences and identifies visually anchored tokens. Notably, the best overall results are obtained with $a=0.5$, validating our proposed method and hypothesis. This indicates that when the visual-anchored score $s=0$, setting $c=1$, not introducing additional reward signals can yield better outcomes.
Figure 5: Performance curves with the change of the noise steps-(a) and the change of parameter aa-(b), We separately present the F1 of AMBER, the hallucination rate of MMHAL, the aACC of HallusionBench, and the average value of the general benchmarks.
@@ 4.4 Ablation Studies
Acc F1 Score Hal $\downarrow$ Easy Hard aAcc SEED MMB LLaVA $M$-Vet
@@ 5 Conclusion
In this study, we propose a novel pereference alignment method, TPO, to mitigation hallucinations in LVLMs. TPO incorporates a self-calibrated visual-anchored reward mechanism that automatically identifies "vision-anchored tokens" and adaptively assigns appropriate rewards to them. By adding noise to the visual input and capturing changes in the generation probability of each token, TPO computes a score indicating each token’s relevance to visual information. Based on the self-calibrated visual-anchored reward, TPO can perform more efficient token-level preference alignment optimization for LVLMs. Experimental results have proved that TPO not only alleviates the hallucination problem but also strengthens the model’s attention to visual input when generating responses.
@@ 6 Limitation
Although our method has achieved outstanding performance in addressing the hallucination problem, the self-calibrated visual-anchored rewards approach we used in this paper can be extended to even broader areas. By altering the way noise is added to images, we can shift from adding noise to the entire image to adding noise to specific key objects. It can enable the model to specifically improve its focus on image information in certain domains, thus having extensive industrial applications. Besides, we believe that the core part of the TPO method, the visual-anchored reward scoring method, possesses strong extensibility. For example, these token-level rewards can also be used to weight the probability distributions in the calculations for other RLHF methods, enhancing the visual attention of multimodal models.
We will continue to expand in this direction, and we believe that the technology we have proposed in this paper has a vast space for further development and application.
@@ A.1 Setup
During the training process, we froze the vision encoder and only trained the LLM.
We also trained the Qwen2-VL model. For our TPO method and the vanilla DPO method, we set the maximum learning rate to 5e-9 for 7B model, 1e-9 for 2B model and trained for 4 epochs.
@@
Our experiments were all conducted on a server equipped with 8 Nivdia A100 GPUs; in specific cases (such as the 13B model), we utilized 32 Nivdia A100 GPUs. For the hyperparameter settings, all hyperparameters are consistent with those of our main experiment. We used ‘Pytorch’ in our code. Moreover, the level of diffusion noise in our model is represented by a formula $\xi=Sigmoid(l_{t})\times(0.5\times 10^{-2}-10^{-5})+10^{-5}$, where $l_{t}$ is a list of 1,000 numbers taken at equal intervals over the interval [6,6][-6,6], and $\epsilon\in N(0,1)$.